A uniform beam in a roof structure has a weight of 180N it is supported in two places X and Y a distance 3.0m apart . A load is placed on the beam a distance of 0.80m from X .The support provided by Y is 220N
What is the value of load?
1
Expert's answer
2020-09-17T14:29:25-0400

Taking moments about Y:


"(3)R_Y-(3\cdot 0.5)W-(0.8)F=0\\(3)220-(3\cdot 0.5)180-(0.8)F=0\\F=487.5\ N"